OpenStudy (anonymous):

Can someone please help me Convert as indicated. 5 lb to grams A. 2100.0 g B. 2270.0 g C. 2325.0 g D. 2240.0 g

OpenStudy (mathstudent55):

Do you know how many grams there are in a lb?

OpenStudy (anonymous):

1 lb = 453.592 g

OpenStudy (anonymous):

I got B but i got it wrong

OpenStudy (anonymous):

so am kind of confuse

OpenStudy (anonymous):

\[453.592*5=2267.96 \approx B.\]

OpenStudy (anonymous):

that's what I got but they mark it wrong don't know why

OpenStudy (anonymous):

that's weird..

OpenStudy (anonymous):

it sure is

OpenStudy (mathstudent55):

\( 1 ~lb \approx 0.454 ~g\) Multiply both sides by 5: \( 5 \times 1 ~lb \approx 5 \times 0.454 ~g\) \( 5 ~lb \approx 2270.0 ~g\) B is correct. It's the closest to the exact number.

OpenStudy (anonymous):

I know but when I submit my test i got it wrong

OpenStudy (anonymous):

can't really help you there :/

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

ask your teacher maybe?

OpenStudy (mathstudent55):

Maybe you should be grading the test instead of having your test graded.

OpenStudy (anonymous):

yeah, you've beat the test! :D

OpenStudy (anonymous):

lol yeap

OpenStudy (anonymous):

am about to just guess any other answer

OpenStudy (anonymous):

kk good idea

OpenStudy (anonymous):

can some one help me with this

OpenStudy (anonymous):

???

OpenStudy (anonymous):

i got it thanks

OpenStudy (anonymous):



OpenStudy (anonymous):

The figures are similar. Find the length of the missing side indicated with an x or y. A. x = 7.2; y = 3.9 B. x = 6.3; y = 3.6 C. x = 6.3; y = 3.9 D. x = 1.24; y = 3.6

OpenStudy (anonymous):

need help with this problem please anyone?

OpenStudy (mathstudent55):

In the future, start a new post for each new problem.

OpenStudy (mathstudent55):

Since the triangles are similar, the lengths of corresponding sides are in proportion. That means the ratios of the lengths of corresponding sides are equal. You need one pair of corresponding sides that will tell you what that ratio is. Looking at the figure, you see that: Side y corresponds to side 8.1. Side 2.8 coprresponds to side x. Side 2.4 corresponds to side 3.4. The only pair of corresponding sides that has given lengths of both sides is the pair 2.4 and 3.4 That tell us the ratio of the lengths of corresponding sides is \( \dfrac{2.4}{3.4} \) We also have the ratios: \( \dfrac{y}{8.1} \) and \( \dfrac{2.8}{x} \) If you set up proportions using the known ratio and each one of the ratios containiong an unknown length, then you'll have two proportions to solve, and you'll be able to find x and y. \( \dfrac{2.4}{3.4} = \dfrac{y}{8.1} \) Solve for y. \( \dfrac{2.4}{3.4} = \dfrac{2.8}{x} \) Solve for x.

OpenStudy (anonymous):

the answer is B